OpenStudy (anonymous):

what is the solution to the equation cos2x+2=sinx for 0degrees

OpenStudy (anonymous):

Use the trig identity: cos2x=1-2(sinx)^2 1-2(sinx)^2-sinx+2=0 -2(sinx)^2-sinx+1=0 (Sinx+1)(2sinx-1)=0 Sinx=-1 or sinx=1/2 X=3pi/2 or x= pi/6 or x=5pi/6